# Tru-POMDP: Task Planning Under Uncertainty via Tree of Hypotheses and Open-Ended POMDPs

### v1

@@ A.2 LLM-Generated Rollout Policy in Online POMDP Planning
Role:
You are helping a robot find the target objects in the environment. The robots state is described by a SceneGraphSimple object, which tracks:
@@
There is also a list of goals, where each goal indicates a goal_object_id. The robot must open the parent areas of goal objects, to make them observable.
@@
Write a C++ function named NextActionExp that determines the next single action the robot should take to work toward completing any remaining goals. After performing the returned action, the function may be called again until all goals are met.
@@
ActionSimple NextActionExp(
const vector<int> &unreached_goals,
@@
1. unreached_goals is a list of object_id describing which objects still not observable.
@@
[…]
@@
[…]
@@
1. To make one object observable, you must open its parent area.
2. If there is no meaningful action to take, you must return ActionSimple() with no arguments (the no-op action).
@@
if (!current_scene_graph.CheckAreaInScene(goal_area) ||
@@ A.3 Kitchen Environments and Area Naming
We select five kitchen environments from RoboCasa nasiriany2024robocasa . The layouts are illustrated in Figure 6. We remove surface-mounted furniture items (e.g., sinks) while retaining furniture with internal storage spaces (e.g., ovens).
(a) One Wall
(b) One Wall with Island
(c) L-Shaped
(d) L-Shaped with Island
(e) Galley
Figure 6: Kitchen Environments from RoboCasa nasiriany2024robocasa
@@
    {tblr}
    width = colspec = Q[1.6cm,c,m] Q[1cm,c,m] X[c,m], hlines, vlines, cells = font= Category & Number Area Names
@@
    {tblr}
    width = colspec = Q[1.6cm,c,m] Q[1cm,c,m] X[c,m], hlines, vlines, cells = font= Category & Number Area Names
@@
    {tblr}
    width = colspec = Q[1.6cm,c,m] Q[1cm,c,m] X[c,m], hlines, vlines, cells = font= Category & Number Area Names
@@
    {tblr}
    width = colspec = Q[1.6cm,c,m] Q[1cm,c,m] X[c,m], hlines, vlines, cells = font= Category & Number Area Names
@@
    {tblr}
    width = colspec = Q[1.6cm,c,m] Q[1cm,c,m] X[c,m], hlines, vlines, cells = font= Category & Number Area Names
@@ A.5 Hyperparameters in Tru-POMDP
{tblr}
width = colspec = Q[1.5cm,c,m] X[c,m] Q[1cm,c,m], hlines, vlines, row1 = font=, cells = font= Parameter & Description Value

### v2

@@ Acknowledgment
This work was supported in part by the National Key R&D Program of China (Grant No. 2024YFB4707600) and the National Natural Science Foundation of China (Grant No. 62303304).
@@ A.2 LLM-Generated Rollout Policy in Online POMDP Planning
You are helping a robot perform an object-rearrangement task in a simulated environment. The robots state is described by a SceneGraphSimple object, which tracks:
@@
4. Which object, if any, is currently being held by the robot.
There is also a list of goals, where each goal indicates a pairing of [goal_area_id, goal_object_id]. The robot must move each specified object into the specified area.
@@
Write a C++ function named NextAction that determines the next single action the robot should take to work toward completing any remaining goals. After performing the returned action, the function may be called again until all goals are met.
@@
ActionSimple NextAction(
const std::vctor<std::array<int, 2>>& unreached_goals,
@@
1. unreached_goals is a list of [area_id, object_id] pairs describing which objects still need to be moved to which areas.
@@
int GetObjectInHand() const;
// Returns the ID of the object currently in the robots hand. -1 if none.
@@
ACTION_OPEN,
ACTION_PICK,
ACTION_PLACE
@@
1. ActionSimple(ActionType::ACTION_OPEN, area_id, -1) tells the robot to open the specified area.
2. ActionSimple(ActionType::ACTION_PICK, area_id, object_id) instructs the robot to pick up an object from an area.
3. ActionSimple(ActionType::ACTION_PLACE, area_id, -1) instructs the robot to place its currently held object into the specified area.
@@
1. Before taking objects from an closed area, you need to open it first. You can only pick one object at the same time. Generate repeated actions is allowed, but do not generate looped actions.
2. You need to find the target object and perform the action in the correct order. You must interact with objects that are observed. You must generate the defined actions.
If there is no meaningful action to take, you must return ActionSimple() with no arguments (the no-op action).
@@
if (!current_scene_graph.CheckAreaInScene(goal_area) $|$
@@ A.3 Kitchen Environments and Area Naming
We select five kitchen environments from RoboCasa Nasiriany et al. (2024). The layouts are illustrated in Figure 6(e). We remove surface-mounted furniture items (e.g., sinks) while retaining furniture with internal storage spaces (e.g., ovens).
((a))
((b))
((c))
((d))
((e))
Figure 6: Kitchen Environments from RoboCasa Nasiriany et al. (2024)
@@
    Category Number Area Names
@@
    Category Number Area Names
@@
    Category Number Area Names
@@
    Category Number Area Names
@@
    Category Number Area Names
@@ A.5 Hyperparameters in Tru-POMDP
Parameter Description Value